how can i multiply 30%(-0.8)
Answer by stanbon(75887)

multiply 30%(-0.8)
----
= (30/100)(-8/10)
---
= (3/10)(-8/10)
----
= -24/100
----
= -0.24
================
Cheers,
Stan H.